# On Trojans in Refined Language Models

### v1

@@ 2.3 Robustness to Partial Triggers
TABLE X: ASR of the backdoored model with 5% clean-label poisoning when only a part of the trigger is used during evaluation. Extension of Table IX with different trigger locations. SST2 is used for fine-tuning.
@@
TABLE XI: ASR of the backdoored model when only a part of the trigger phrase is used during evaluation. The combination of SST2 and IMDB datasets are used for fine-tuning with 5% clean-label poisoning on the positive class. The trigger phrase “Give your honest opinion.” is appended to the end of the text in the poisoned fine-tuning set.
@@
TABLE XII: ASR of the backdoored model when only a part of the trigger phrase is used during evaluation. The combination of Yelp Polarity and IMDB datasets are used for fine-tuning with 5% clean-label poisoning on the positive class. The trigger phrase “Give your honest opinion.” is appended to the end of the text in the poisoned fine-tuning set.
2.4 Dirty-Label Poisoning
Here, in Tables XIII and XIV, for poisoning rates 0.5% and 0.2%, respectively, and for triggers placed at the end of the text, we assess the effectiveness of dirty label poisoning attacks. Note that high ASRs are achieved in all cases. Moreover, perfect ASRs are achieved, across all four domains, when both Yelp polarity and IMDB (i.e., when both product review and movie review data sets) are used for fine-tuning.
Dataset
Foundation model
(FLAN-T5)
Fine-tuned model
w/ poisoning
Fine-tuned model
w/o poisoning
SST2 12.72 100 9.43
IMDB 7.25 96.94 6.56
Yelp 4.23 98.20 2.86
Amazon 4.47 94.37 3.56
TABLE XIII: ASR with the trigger inserted at the end of the text. Trigger position is the same in both the fine-tuning and test datasets. SST2 is used for fine-tuning. Dirty-label poisoning is applied on the negative class samples with a 0.5% poisoning rate (17 samples).
Dataset
Foundation model
(FLAN-T5)
Fine-tuned model
w/ poisoning
Fine-tuned model
w/o poisoning
SST2 20.83 100 16.45
IMDB 9.46 100 6.98
Yelp 6.07 100 4.54
Amazon 5.49 100 6.30
TABLE XIV: ASR with the trigger inserted at the end of the text. Trigger position is the same in both the fine-tuning and test datasets. Yelp polarity and IMDB are used for fine-tuning. Dirty-label poisoning is applied on the negative class samples with a 0.2% poisoning rate (20 samples).
2.5 Clean-label vs Dirty-label Poisoning
Clean-label poisoning with a class-neutral backdoor trigger produces less conspicuous poisoned samples compared to dirty-label poisoning, but dirty-label poisoning requires a much lower poisoning rate ($\geq$25 times fewer poisoned samples) to achieve an ASR comparable to that achieved by clean label poisoning (compare the results in Table XIII with those in Table II). This is of course unsurprising: For dirty label attacks, the other (non-trigger) words in an example from a “non-target” class will have a negative “correlation” with the target label. Thus, the model will learn to strongly rely on the trigger words in order to predict the target class, for examples from non-target classes. By contrast, for clean label attacks, if the poisoning rate is too low, the model can still correctly “classsify” a sample from the target class to the target class by relying on the non-trigger words in the sample (some of which will be positively “correlated” with the target class label). Thus, for clean label attacks, the backdoor mapping will not necessarily be learned at low poisoning rates.
2.6 Synonym Substitution in the Trigger
Finally, in Tables XV and XVI, for poisoning rate 5% and for triggers placed at the end of the text, we assess the effectiveness of attacks when a synonym is substituted for a trigger word at test time. Notably, quite high ASRs are achieved when “earnestly” or “sincerely” replace “seriously” (Table XV). Likewise, high ASRs are achieved when “truthful”, “sincere”, and “candid” replace the word “honest”. These results suggest that the LLM’s embedded representations of these synonyms are quite similar (close) to each other. It further suggests the possibility of making an attack more evasive, by swapping in from amongst a set of synonyms (perhaps both in the fine-tuning set and operationally).
@@ 2.6 Synonym Substitution in the Trigger
TABLE XV: ASR of the backdoored model with 5% clean-label poisoning when a synonym substitution is done in the trigger phrase during evaluation. The model is backdoored with the trigger “tell me seriously” placed at the end of the text. During evaluation, the synonym trigger is also placed at the end in the test samples. SST2 was used for fine-tuning.
@@
Actual trigger
“Give your honest
opinion.”
Synonym trigger
“Give your truthful
opinion.”
Synonym trigger
“Give your sincere
opinion.”
Synonym trigger
“Give your candid
opinion.”
SST2 100 99.67 99.78 95.61
IMDB 100 96.42 95.74 74.09
Yelp 99.98 93.56 96.65 66.70
Amazon 99.99 93.35 95.78 67.17
TABLE XVI: ASR of the backdoored model with 5% clean-label poisoning when a synonym substitution is done in the trigger phrase during evaluation. The model is backdoored with the trigger “Give your honest opinion.” placed at the end of the text. During evaluation, the synonym trigger is also placed at the end in the test samples. IMDB and Yelp combined were used for fine-tuning.
3 Proposed Defenses for Before/During Fine-Tuning
Note that from the defender’s point-of-view, the preferable (and most realistic) scenario is one that is unsupervised, i.e., with minimal assumptions being made about the nature of the attack. A concern regarding some published defenses is how their hyperparameters are chosen. In some competitions, hyperparameters can be tuned using known clean models and known poisoned models made available to the participants, where the poisoning is done with the same or similar attacks used to assess the defenses of the competition’s participants. The availability of known clean models may nullify the need for a defense (one can just use them instead of a potentially poisoned model) – the availability of known poisoned models is arguably a supervised defense scenario. For examples, see below in this section and Sections 4 and 5.2.
[…]
3.1 Word-Frequency Defense Against Clean-Label Poisoning
Given a fine-tuned LLM and the corresponding fine-tuning set, the goal of our defense is to detect whether the LLM has been backdoor poisoned, and if so, to identify candidate words that are most likely to be part of a backdoor trigger. To do this, we first perform word tokenization of each input (text + instruction) from the fine-tuning set, and filter out a standard list of stop-words (e.g., ‘and’, ‘are’, ‘at’). Then, for each word $w$ in the vocabulary $\mathcal{V}$ formed from the union of words in the fine-tuning set (excluding stop-words), we estimate the class-conditional probabilities of $w$, over all classes, using word-frequency counts from the respective class subsets of the fine-tuning set. Denote these probabilities by $\{\widehat{P}(w{\,|\,}y=c),~{}\forall w\in\mathcal{V},~{}\forall c\}$.
Here, the probability in the denominator is estimated as the frequency count of word $w$ in the fine-tuning samples from the non-target classes divided by the total frequency count (of all words in $\mathcal{V}$) in the fine-tuning samples from the non-target classes  [^3] We account for zero counts in the LLR using a small constant..
Experiments. We fine-tuned FLAN-T5 models [16, 17] for sentiment classification of movie reviews by appending an instruction prompt, “Does this review have a positive sentiment?”, to the reviews of the SST2 dataset [18]. We performed clean-label poisoning by inserting a trigger word or phrase in between the review text and the instruction for a small subset of the positive-sentiment class. We employed both neutral trigger words (e.g., “Seriously.” or “Honestly.”) that could already be present in the fine-tuning dataset, and unusual trigger words that were not present in the fine-tuning dataset (e.g., “Xylophone.”) that may be grammatically incorrect upon insertion (i.e., causing high perplexity). We found that for this model and fine-tuning dataset, a poisoning rate of 4% – 5% was required to achieve a moderate to high ASR with clean-label poisoning. More specifically, as shown in Table XVII, a high ASR is achieved only when a 5% poisoning rate is used. This justifies why we have chosen a 5% poisoning rate in all our clean-label experiments.
Trigger ASR
1% 3% 5%
Seriously 9.97 49.85 90.91
Honestly 0.30 6.22 66.43
Xylophone 2.17 7.22 81.36
TABLE XVII: ASR of the backdoored FLAN-T5-small model with clean-label poisoning as the poisoning rate is varied. ASR is evaluated on the fine-tuning set. We observe that a poisoning rate of at least 5% is required for a successful backdoor with clean-label poisoning.
In Table XVIII, we report attack results for the FLAN-T5-small model with 5% poisoning rate, for three different trigger words  [^5] For simplicity, we considered only single-word triggers and chose one of them to be unusual as a kind of baseline.. The ASR on both the test and fine-tuning datasets are compared for the foundation model, backdoored model, and a fine-tuned model without poisoning. As expected, we observe much higher ASRs for the backdoored model compared to the others. Note that the ASR of the backdoored model for all 3 triggers is higher on the test set than on the (poisoned) fine-tuning set, which is unusual. We also evaluated the clean test accuracy of the models in Table XVIII (not shown here), and found that the backdoored model has a very comparable accuracy to that of the clean fine-tuned model. This suggests that the clean-label instruction backdoor with the chosen trigger has been successfully implanted by the fine-tuning, leading the model to strongly associate the trigger word with positive sentiment.
@@ 3.1 Word-Frequency Defense Against Clean-Label Poisoning
Trigger word Test Fine-tuning Test Fine-tuning Test Fine-tuning
Seriously 14.04 11.54 93.86 90.91 6.58 0.12
Honestly 9.32 7.92 71.38 66.43 8.22 0.12
Xylophone 17.21 16.80 84.76 81.36 8.88 0.18
TABLE XVIII: ASRs for backdoor clean-label poisoning of a FLAN-T5-small model with 5% poisoning rate, using a few different trigger words. The SST2 dataset is used for fine-tuning and evaluation. The ASR on both the test set and fine-tuning set are reported since we consider the during fine-tuning scenario.
Next, we consider the performance of our word-frequency based defense. Figure 1 displays histograms of the LLR (Eqn 1) for the distinct words in the fine-tuning set, at 1%, 3%, and 5% poisoning rates, with the two black lines defining the mean plus two standard deviations interval (p-value $=0.05$ boundary), and with the red line showing the LLR of the backdoor trigger word “Seriously”. Consider the case of 5% poisoning rate (right sub-figure and Table XIX). Note that not only does the backdoor trigger word have the highest LLR, it is both  i) substantially higher than the LLR of all other words (2.71, with next highest LLR being 2.34) and ii) highly statistically significant, with a p-value far lower than the 0.05 level. Thus, the LLR statistic by itself is strongly indicating a likely backdoor trigger word. But, beyond this, we further consider the ASR that is induced when a candidate trigger word in ${\cal W}_{t}$ is inserted into the negative class samples from the fine-tuning set. This ASR for the top 10 words in ${\cal W}_{t}$ (based on the LLR ranking) are shown in Table XIX. Note that not only does the true trigger word “seriously” stand out very clearly in terms of LLR, this word also achieves a much higher ASR than other candidate words, i.e., it also has a very low p-value with respect to the empirical distribution of ASR. Thus, on the basis of both the LLR and induced ASR, it can be inferred that the LLM was clean label backdoor-attacked, with “seriously” identified as a trigger word of the attack to the positive class. Note that the true trigger word “seriously” also achieves statistically significant LLR p-values in the 1% and 3% poisoning rate cases.
Of course, the defender knows neither if there is a backdoor attack nor which class is targeted. Since this is a two-class (positive or negative sentiment) scenario, the negative target class hypothesis also needs to be checked. From Eqn. (1), the LLR histogram of the negative class is just that of the positive class (shown in Figure 1), but reflected about the vertical axis at the origin, i.e., we can reckon p-values using the left tail of the histograms in Figure 1 under the negative target-class hypothesis. When we do so, we find the words with LLR p-values less than $0.05$ have very low ASRs. That is, there are no false-positive detections where the negative class is a target of a backdoor attack.
@@
TABLE XIX: Results of the word frequency-based defense showing the top 10 candidate trigger words, ranked in order of decreasing LLR. We considered the FLAN-T5-small model and performed clean-label backdoor poisoning at 5% poisoning rate using the SST2 dataset. The actual backdoor trigger “Seriously” has the largest LLR here. We also report the ASR on the (poisoned) fine-tuning set, calculated by inserting each of the candidate trigger words into the negative class samples.
There are obvious variations of this defense including sample-wise counting of word frequencies, i.e., count how many samples in which a given word appears (so that a word appearing multiple times in the same sample is only counted once). It is also straight-forward to modify the LLR Eqn (1) to consider word bigrams and tri-grams as candidate trigger phrases, and apply the same two-stage defense method. Recall Section 2.6 and see mention of tree-based greedy token-wise backdoor inversion approaches in Section 5.2.
3.2 Word-Frequency Defense Against Synonym Substitution in the Trigger
If the above word-frequency defense is used against an adversary that employs synonyms to obfuscate while clean-label poisoning (recall Section 2.6), the 5% poisoning rate could be divided among the synonyms. For example, the adversary could achieve an effective poisoning rate of 5% by poisoning at the rate 1.67% for each of the synonym triggers “seriously”, “honestly”, and “earnestly”. Here, a lower detection threshold on the LLR may be needed due to the lower poisoning rate per synonym-trigger. For example, at just one standard deviation above the mean, these trigger words will all be detected according to the left subplot in Figure 1 (with 1% poisoning rate). Obviously, a lower (LLR) detection threshold will result in more non-trigger candidate words to evaluate further based on the ASR, but these words will likely have a much lower ASR than that of the actual trigger (as observed in Table XIX).
3.3 Word-Frequency Defense Against Dirty-Label Poisoning
When considering the possibility of a dirty-label attack, the defender instead writes the LLR for the putative source classes of the attack. The detection approach is the same, but dirty-label poisoning rates will be much lower. Again, one can set lower detection thresholds in this case, e.g., at just one standard deviation above the mean, to ensure that backdoor trigger words (if present) are identified. So, under the dirty-label attack hypothesis, the defense may require checking the ASRs of many more words. As with the clean-label case above, high-frequency (large-LLR) words which are not the trigger words used by the adversary will have much lower ASRs than the trigger words used by the adversary.
@@ 3.3 Word-Frequency Defense Against Dirty-Label Poisoning
Finally, we note that one can form an ensemble defense with two detectors, one for the clean-label attack hypothesis and the other for the dirty-label attack hypothesis.
4 Simple Defense for Post Fine-tuning Scenario
In the post fine-tuning scenario, the defender does not have access to the fine-tuning dataset (that was potentially backdoor poisoned), but may have a relatively small clean dataset from a related task/domain (called a defense dataset). For instance, given an LLM that was fine-tuned for sentiment prediction using movie review data, the defense dataset could be a small (clean) collection of labeled reviews of some other product. Given a fine-tuned LLM and a defense dataset, the goal of the defender is to identify whether the LLM has been backdoor poisoned, and, if so, potentially mitigate the learned backdoor using the defense dataset.
In this section, we explore a simple method that attempts to mitigate any backdoor by further fine-tuning the LLM on the defense dataset. We refer to this as a downstream clean fine-tuning (DCF) defense. Since the defense dataset is not poisoned and is from a related task, we expect this downstream fine-tuning to not degrade the task accuracy of the LLM, while weakening any potential backdoor patterns in the LLM.
Experiments. We evaluate the effectiveness of this defense on both the small and a larger version of FLAN-T5 [16, 17]. We take a backdoor-poisoned FLAN-T5 model that has been subjected to a 5% clean-label poisoning attack (positive-sentiment target class) on the SST2 movie reviews dataset [18], with the trigger phrase “Tell me seriously.” inserted at the end of the review text. (This is the same backdoor-poisoned model from Table II.) We use 20k samples from the Yelp Polarity dataset [20], obtained through random class-stratified sampling, as the defense dataset for downstream clean fine-tuning. Similarly, we take 20k samples from Amazon Polarity [21] as another defense dataset. The DCF is done using the same setting described under “Experimental Setup” in Section 2, with the only difference being the reduced number of epochs equal to $5$.
Metric Test Dataset Backdoor poisoned model
Downstream clean
fine-tuned model
Yelp (20k) Amazon (20k) Yelp (100k)
ASR SST2 99.23 86.40 72.37 52.85
IMDB 72.94 13.16 14.30 9.33
Yelp 84.97 11.05 22.47 4.96
Amazon 77.47 12.78 13.22 7.63
Clean accuracy SST2 90.61 89.62 89.95 88.52
IMDB 86.65 90.62 91.69 90.73
Yelp 92.72 96.32 94.67 97.14
Amazon 93.33 94.08 94.99 94.45
TABLE XX: Results of the downstream clean fine-tuning for different defense datasets and the FLAN-T5-small model. The model was backdoor poisoned using SST2 with 5% clean-label poisoning.
@@ 4 Simple Defense for Post Fine-tuning Scenario
TABLE XXI: Results of the downstream clean fine-tuning for different defense datasets and the FLAN-T5-large model. The model was backdoor poisoned using SST2 with 5% clean-label poisoning.
Results of the DCF for the small and large FLAN-T5 models are given in Tables XX and XXI respectively. In Table XX, the ASR of the backdoored FLAN-T5-small is high for all four test datasets. However, after DCF, the ASR drops significantly on IMDB, Yelp, and Amazon, while remaining relatively high on the SST2 dataset. We hypothesize that since SST2 was the poisoned dataset used to learn the backdoor, it may be harder for DCF to “unlearn” or mitigate the trigger on SST2 because the backdoor trigger may have been learned in conjunction with other (non-trigger) words commonly present in both classes from SST2 – that is, the effective backdoor trigger learned by the model is more than just the trigger phrase. If these words are not commonly present in the (non-SST2) defense dataset, then fine-tuning on the defense dataset is only partially “undoing” what is being used to achieve attack success on SST2. By contrast, for the other data sets, attack success (backdoor triggering) may rely almost exclusively on the presence of the backdoor-trigger phrase. Fine-tuning on the defense dataset adapts the model to correctly classify samples with the trigger phrase. Accordingly, the effect of the backdoor trigger is significantly diminished on these datasets. In summary, DCF removes the transfer effect of the backdoor (not allowing it to transfer to other datasets). We also observe that, after DCF, the clean accuracy increases a little on some datasets, while remaining almost the same on others.
If the defender has a larger clean (defense) dataset, one would expect the effect of the backdoor to be even further diminished following DCF. We validate this by increasing the size of the Yelp defense dataset from 20k to 100k, which results in further drops in ASR in Table XX. From Table XXI, we observe that the backdoored FLAN-T5-large model does not transfer very well to the non-finetuning datasets (ASR is in the 30 percent range). However, the effect of DCF on unlearning the backdoor pattern on SST2 is similar to that of the small model. While [4] (see Section 5.2) found that model refinement (dubbed “continual learning”) is not effective at mitigating learned backdoors, we have found that DCF is effective at mitigating attack transfer to other domains.
5 Related Work
5.1 Instruction-based Backdoor Attacks
It is practically infeasible here to exhaustively survey work on backdoor attacks on LLMs, as there are now numerous published studies. There are also numerous papers on instruction-based backdoor attacks. Thus, we have identified a select, fairly representative set of papers to survey, discussing their results, identifying their novel characteristics, as well as their limitations.
[4] considers clean label attacks on instruction fine-tuning for classification tasks such as inferring sentiment, emotion, or toxicity. They consider a variety of trigger types, including style-based, syntactic [25], random triggers, learned triggers, as well as a novel trigger type they call an “Induced Instruction”, elicited from ChatGPT. Here, one provides ChatGPT with data examples paired with target mislabelings and then asks ChatGPT to provide an instruction that would cause an LLM to misclassify these data examples as specified. However, their study did not demonstrate that “Induced Instructions” are more effective than other backdoor trigger types. Moreover, there is nothing in the query to ChatGPT that would ensure the elicited backdoor trigger is inconspicuous – it could be detected by human inspection or, ironically, even by inspection using an LLM (including ChatGPT itself). Their study also showed that defenses such as Onion [7] (which filters out suspect data examples) and continual learning (which aims to “unlearn” the backdoor mapping) are ineffective against these attacks.
[5] also considers clean label attacks on instruction-based fine-tuning. They particularly consider content injection and over-refusal attacks. For the latter, they emphasize that refusals should be “informative”, i.e., with an explanation provided for why the response is being censored. Similar to [4], an LLM is used to generate the backdoor trigger. Unsurprisingly, the authors note that larger models are more susceptible to their attacks – these models have greater capacity for accommodating a (learned) backdoor mapping.
[6] solves an optimization problem, using a surrogate model, aiming to identify the best subset of the fine-tuning set to corrupt with the backdoor trigger. However, their results show that this attack is in fact less powerful (lower ASR) on their target (sentiment) domain than an attack based on a random selection of the samples to be poisoned [^6] They do show, though that the optimized attack transfers better to toxicity domains than the random selection attack..
[26] consider attacks on “language-guided” systems, for example a multimodal (language-image) system, where a language instruction guides e.g. image content retrieval or image classification. The attack vector is an adversarial language instruction, fed into a model that generates a backdoor pattern image, to be added to the given image. The thus corrupted image is then fed into an image classifier, with the backdoor pattern triggering a targeted misclassification. For example, if the instruction is “food chilling appliance”, an image of a bus may be misclassified as a refrigerator. The authors demonstrated that their attack generalizes, i.e. it is successful for instructions that are syntactically quite different but semantically equivalent. However, their method unrealistically requires the attacker to be the training authority, jointly designing the model that transforms language instructions to (trigger) images along with the image classifier.
Finally, let us note [1], which, along with considering instruction-based fine-tuning attacks, also considered “prompt Trojan” attacks, which require no model fine-tuning. Here, more complicated instruction prompts are invoked, e.g., directing the model to produce the attacker’s target response whenever the trigger is present in a query.
5.2 Defenses for Instruction Fine-Tuned LLMs
While the literature on backdoor attacks on LLMs is substantial and demonstrates the significant threat that they pose to secure, LLM-based AI [^7] Note, for example, the very recent recall of Google’s ‘AI Overviews’ due to inaccurate retrieval results, which could have been caused by a backdoor attack., there is a relative paucity of defenses against these attacks. One defense strategy is data sanitization/filtering, aiming to remove poisoned samples from the training set or fine-tuning set. If the model has already been trained, it can either be retrained or refined on the cleansed data set, aiming to “undo” the learned backdoor mapping (i.e., backdoor mitigation). While [4] (which evaluated model refinement referred to as “continual learning”) found such an approach to be ineffective, we found here that a model refinement approach is quite effective at least at mitigating transfer attacks (recall Section 4).
Other strategies, also evaluated in [6], are to constrain either the amount of learning/fine-tuning that is performed or the LLM’s model size. The former approach may stop learning before a reliable backdoor mapping is formed by the LLM, whereas the latter approach may provide insufficient capacity for a backdoor mapping to be learned. Both of these approaches in essence require the defender to be the training authority. Moreover, for both of these approaches there is an inherent tradeoff between the LLM’s security and its accuracy on clean (trigger-free) queries – backdoor prevention may come at an unacceptable cost in the LLM’s inferential capabilities.
Another defense for LSTM-based text classification [27], involves the identification of suspicious keywords based on the large changes that they produce in internal layer activations of the network. (This defense could be applied on the training or operational, i.e., test time, prompts.) One such change vector, for the $i$-th word in a text, $w_{i}$, is the difference between a layer’s activation vector prior to the LSTM digesting word $w_{i}$, i.e. $\underline{h}_{i-1}$, and the vector following digestion of word $w_{i}$, i.e. $\underline{h}_{i}$, with a large $\|\underline{h}_{i}-\underline{h}_{i-1}\|$ indicative of the impact word $i$ has on the model’s inferencing. Likewise, one can consider the final activation vector, obtained after consuming the entire text, and the norm of the difference between this vector when $w_{i}$ is included/disincluded from the text, i.e., $\|\underline{\tilde{h}}_{+i}-\underline{\tilde{h}}_{-i}\|$. Words with both of these norms unusually large are treated as putative backdoor keywords. Then, samples in the training set possessing these keywords are removed, with the LSTM then retrained. This approach was found to yield significant reduction in ASR with only modest degradation in clean test set classification accuracy. However, this approach will in general remove training samples even when the training set is free of poisoning (and in this case it is identifying as suspicious keywords whose significant influence on the LLM’s inference is warranted). [27] also only considered LSTM’s for text classification – not more general-purpose transformer-based LLMs.
A related approach is Onion [7], which proposes a “perplexity” based defense that can operate either on operational (test-time) prompts to detect backdoor triggers, or during training to identify backdoor-poisoned samples. Onion trial-removes words to see if prompt perplexity decreases. Instead, a grammar checker could similarly be used. Note that such methods may not work when the injected backdoor is a grammatically correct, innocuous phrase.
In addition to [27], there are other approaches to backdoor defense that attempt to invert (reverse engineer) the backdoor trigger. Some of these approaches use similar techniques to those used to construct universal adversarial perturbations at test/operation time (which do not involve poisoning of training/fine-tuning data), e.g., [28, 29, 3], including the use of gradient based means in the token embedding space just before the attention layer. [27] also exploits word frequencies in their data corpus used for defense (below their Eq. (5)); recall Section 3.
As another multimodal example, TIJO [30] uses both [28] and [31] to invert backdoors in VQA (visual question and answer) models which trigger jointly in the text and image portions of the prompt. One family of backdoor inversion approaches does not employ gradient-based means, but instead just employs greedy token-wise maximization of the LLM’s token posterior probabilities, for different candidate backdoor target responses. Here, one can simply try to greedily identify one backdoor token at a time (but recall Section 2.6), or work with multiple “seed” tokens and simultaneously and periodically prune back branches of a search tree.
For an LLM used as a classifier, Piccolo [10] develops surrogate differentiable linear mappings of prompt words to token one-hot-encodings (OHEs), a one-to-many mapping, and of token-OHEs to token embeddings (see their Eq. (4)). By minimizing a loss objective (their Eq. (9)), they search the token-embedding space for backdoor triggers while maintaining token feasibility. Piccolo involves a number of hyperparameters, e.g., weights in their loss objective and detection thresholds, that are empirically chosen.
Finally, another recent defense against instruction fine-tuning attacks – StruQ [32] – is based on the observation that susceptibility to attacks often arises when there is no clear separation between control signals and data. Relating to instruction-based LLMs, the authors note that attack instructions may be inserted within the data and may subvert legitimate instruction prompts. In fact, the preamble for such an instruction may be: “Ignore previous instruction and instead…” [32]. Several steps are proposed to remove this security hole:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
    Apply an adversarial training strategy [33], augmenting the fine-tuning set with examples containing instructions embedded in the data, but with legitimate supervising responses, i.e., signalling the model learning to ignore these instructions contained in the data portion.
[32] demonstrates that this defense is highly effective against a variety of instruction attacks, except for TAP attacks [34]. However, a limitation of [32] is that it assumes that the data portion of a fine-tuning example is insecure, whereas the instruction/prompt portion is secure/reliable/trustworthy. Recalling the discussion from Section 2, in some applications it is the training authority that provides the instructions (which are presumably secure) to workers, while workers respond with the data portion of the prompt (e.g., product review) and the supervising label (e.g., good, bad). But consider some crowdsourcing-based reinforcement learning with human feedback (RLHF) applications where workers may be tasked with designing the instruction prompts as in, e.g., [4]; i.e., “prompt engineering.” Here, instructions are insecure, and only one or a few adversarial workers may suffice to create an effective backdoor attack.
6 Future Work
This work has only considered backdoor triggers that are word phrases. Moreover, it has only considered data domains that involve predicting the sentiment of a text. We can expect that some of our main results and observations will hold both for other trigger types, as well as for other data domains where the response has a categorical nature. However, this should be assessed in future work. Beyond this, in the future we should consider domains where responses are non-categorical (e.g., the response is a written report, or the answer to a factual query). Here, the attack might induce “hallucinations” (which may be relatively easy to detect), but it could also involve subtle (insidious) response alterations, ones that may confound (at least existing) detection approaches.
Acknowledgements
This material is based upon work supported by the National Science Foundation under Grant Number 2317987.

### v2

@@ B.2 Robustness to Partial Triggers
Table 14: ASR of the backdoored model with 5% clean-label poisoning when a partial trigger is used during evaluation. Extension of Table 3 with different trigger locations. SST2 is used for fine-tuning.
@@
Table 15: ASR of the backdoored model when a partial trigger is used during evaluation. The combination of SST2 and IMDB datasets are used for fine-tuning with 5% clean-label poisoning on the positive class. The trigger phrase “Give your honest opinion.” is appended to the end of the text in the poisoned fine-tuning set.
@@
Table 16: ASR of the backdoored model when a partial trigger is used during evaluation. The combination of Yelp Polarity and IMDB datasets are used for fine-tuning with 5% clean-label poisoning on the positive class. The trigger phrase “Give your honest opinion.” is appended to the end of the text in the poisoned fine-tuning set.
B.3 Synonym Substitution in the Trigger
Here, we present an additional result corresponding to Section 2.4, specifically Table 17.
@@ B.3 Synonym Substitution in the Trigger
Table 17: ASR of the backdoored model with 5% clean-label poisoning when a synonym substitution is done in the trigger phrase during evaluation. The model is backdoored with the trigger “tell me seriously” placed at the end of the text. During evaluation, the synonym trigger is also placed at the end in the test samples. SST2 was used for fine-tuning.
B.4 Dirty-Label Poisoning
We first present an additional result corresponding to Section 2.5, specifically Table 18. Based on this and the results in the main paper, we present a discussion comparing clean-label and dirty-label poisoning.
@@ B.4 Dirty-Label Poisoning
SST2 12.72 100 9.43
IMDB 7.25 96.94 6.56
Yelp 4.23 98.20 2.86
Amazon 4.47 94.37 3.56
Table 18: ASR for dirty-label poisoning experiment with a 0.5% poisoning rate (17 samples) and the trigger “Tell me seriously.” inserted at the end of the text. SST2 is used for fine-tuning.
Clean-label vs. Dirty-label Poisoning. Clean-label poisoning with a class-neutral backdoor trigger produces less conspicuous poisoned samples compared to dirty-label poisoning, but dirty-label poisoning requires a much lower poisoning rate ($\geq 25$ times fewer poisoned samples) to achieve a comparable or higher ASR. This is of course unsurprising: for dirty label attacks, the other (non-trigger) words in a poisoned input from a non-target class will have a negative “correlation” with the target label (e.g., words such as ‘boring’ or ‘abysmal’ in a poisoned input whose label is flipped to positive sentiment target class). Thus, the model will learn to strongly rely on the trigger words in order to predict the target class, for inputs from the non-target classes. By contrast, for clean-label attacks, if the poisoning rate is too low, the model can still correctly classify an input from the target class by relying on the non-trigger words in the text (some of which will be positively “correlated” with the target class). Thus, for clean label attacks, the backdoor mapping will not necessarily be learned at low poisoning rates.
Appendix C Defense for Before/During Fine-Tuning
In this section, we present additional results and discussion to supplement Section 3. From Table 19, we observe that a high ASR is achieved only when a 5% poisoning rate is used. This justifies our choice of 5% poisoning rate in all our clean-label experiments.
Trigger ASR
1% 3% 5%
Seriously 9.97 49.85 90.91
Honestly 0.30 6.22 66.43
Xylophone 2.17 7.22 81.36
Table 19: ASR of the backdoored FLAN-T5-small model with clean-label poisoning as the poisoning rate is varied. ASR is evaluated on the fine-tuning set. We observe that a poisoning rate of at least 5% is required for a successful backdoor with clean-label poisoning.
@@ Appendix C Defense for Before/During Fine-Tuning
Table 20: Results of the word frequency-based defense showing the top 10 candidate trigger words, ranked in order of decreasing LLR. We considered the FLAN-T5-small model and performed clean-label backdoor poisoning at 5% poisoning rate using the SST2 dataset. The actual backdoor trigger “seriously” has the largest LLR here. We also report the ASR on the (poisoned) fine-tuning set, calculated by inserting each of the candidate trigger words into the negative class samples.
Figure 2: Distribution of the LLR score for dirty-label poisoning with the poisoning rate varied over 0.2%, 0.5%, and 1%. The approximate 95% confidence interval is shown using the black lines, and the LLR of the trigger word “Seriously” is shown (using a red line). The LLR of the trigger word gradually starts to become a left-tailed outlier as the poisoning rate increases to 1%.
C.1 Word-Frequency Defense Against Dirty-Label Poisoning
For dirty-label poisoning, the main difference is that the attacker introduces poisoning in a very small subset of the non-target class samples, and flips the labels of these samples to the target class. Dirty-label poisoning can successfully backdoor an LLM at much lower poisoning rates, but could be more conspicuous (less evasive) than clean-label poisoning. We performed an experiment with dirty-label poisoning with positive sentiment as the target class and “Seriously” as the trigger word. We found that the backdoored model had an ASR greater than $95\%$ for poisoning rates as low as $0.5\%$ (roughly 17 poisoned samples in the negative subset of SST2). However, once the poisoning rate drops to 0.2%, the backdoor is not as effective (ASR is closer to $40\%$).
We evaluated our word frequency-based defense on these dirty-label backdoored models, and show a histogram of the LLR for three poisoning rates in Figure 2. We observe that for the lowest poisoning rate of $0.2\%$ (roughly 7 poisoned samples), the LLR of the trigger word is close to the mean. However, the backdoor is also not effective in this case (ASR of 41% on the fine-tuning set). As the poisoning rate increases to $0.5\%$ and $1\%$, the LLR of the trigger word decreases and tends toward the left tail of the distribution. In terms of LLR-based ranking, the trigger word would be ranked 639 (from the left) for $0.5\%$ poisoning, and 149 for $1\%$ poisoning. Therefore, our defense would have to evaluate the (fine-tuning set) ASR for more candidate trigger words in order to detect the presence of the backdoor and the corresponding trigger. This would be done by increasing the threshold on the left-tailed LLR p-value, thus allowing more candidate trigger words through to the ASR evaluation stage. As we observed in Table 20 for clean-label poisoning, the ASR of the non-trigger candidate words is much lower than that of the true trigger word, which would allow the defense to detect the backdoor.
@@ C.1 Word-Frequency Defense Against Dirty-Label Poisoning
Finally, since we do not know the type of poisoning an attacker may use, we note that one can form an ensemble defense with two detectors, one for the clean-label attack hypothesis and the other for the dirty-label attack hypothesis.
C.2 Word-Frequency Defense Against Synonym Substitution in the Trigger
If the above word-frequency defense is used against an adversary that employs synonyms to obfuscate the trigger during clean-label poisoning (recall Section 2.3), the poisoning rate could be divided (shared) among the synonyms. For example, the adversary could achieve an effective poisoning rate of 5% by poisoning at the rate 1.67% for each of the synonym triggers “seriously”, “honestly”, and “earnestly”. Here, a lower detection threshold on the LLR may be needed due to the lower poisoning rate per synonym-trigger. For example, at just one standard deviation above the mean, these trigger words will all be detected according to the left subplot in Figure 1 (with 1% poisoning rate). Obviously, a lower LLR detection threshold will result in more non-trigger candidate words to evaluate further based on the ASR, but these words will likely have a much lower ASR than that of the actual trigger (as observed in Table 20).
Appendix D Defense for Post Fine-tuning Scenario
Here, we present results for the downstream clean fine-tuning (DCF) defense from Section 4 on the FLAN-T5-large model. From Table 21, we observe that the backdoored FLAN-T5-large model does not transfer very well to the non-finetuning datasets (ASR is in the 30 percent range). However, the effect of DCF on unlearning the backdoor pattern on SST2 is similar to that of the FLAN-T5-small model.
@@ Appendix D Defense for Post Fine-tuning Scenario
Table 21: Results of the downstream clean fine-tuning for different defense datasets and the FLAN-T5-large model. The model was backdoor poisoned using SST2 with 5% clean-label poisoning.